Three years ago, Milcah is 3 times older than Joshua. 4 years from now, Milcah is 8 years older than Joshua. Find Joshua's and Milcah's age.
Answer by ankor@dixie-net.com(22740)

Three years ago, Milcah is 3 times older than Joshua. 4 years from now, Milcah is 8 years older than Joshua. Find Joshua's and Milcah's age.
:
M = M's age now
J = J's age now
:
Write what it says:
"Three years ago, Milcah is 3 times older than Joshua."
(M-3) = 3(J-3)
M - 3 = 3J - 9; multiplied what's inside the brackets
M = 3J - 9 + 3
M = (3J-6)
:
"4 years from now, Milcah is 8 years older than Joshua."
M+4 = (J+4) + 8
M + 4 = J + 12
M = J + 12 - 4
M = J + 8
:
Find Joshua's and Milcah's age.
Substitute (3J-6) for M in the above equation:
3J - 6 = J + 8
3J - J = 8 + 6
2J = 14
J = 14/2
J = 7 yrs; J's age now
:
Find M:
M = J + 8
M = 7 + 8
M = 15 yrs; M's age now
:
:
Check solution using the statement:
"Three years ago, Milcah is 3 times older than Joshua."
(15-3) = 3(7-3)
12 = 3(4); confirms our solution